1/2

N. N. 102,

God save the King

God save our gracious King, Long live our noble King, God save The King! Send him victorious, Happy and glorious, Long to reign over us, God save The King.

Thy choicest gifts in store On him be pleased to pour, Long may he reign; May he defend our laws, And ever give us cause To sing with heart and voice, God save The King!

God bless our native land, May heaven’s protective hand Still guard our shore; May peace her power extend, Foe be transformed to friend, And Britain’s power depend On war no more.

…

2/2

May just and righteous laws Uphold the public cause, And bless our isle. Home of the brave and free, Fair land and liberty, We pray that still on thee Kind heaven may smile.

And not this land alone-But be thy mercies known From shore to shore. Lord, make the nations see That men should brothers be, And from one family The wide world o’er.

Verkinto de tiu ĉi Angla poemo estas N. N. 102.

Arg-1097-2212 (2014-11-04 14:51:11)

Tiu ĉi origina teksto de la angla nacia himno troviĝas en http: // en. wikipedia. org/ wiki/ God_Save_the_King# . La nuna oficiala angla nacia himno havas nur tri strofojn.